OpenStudy (anonymous):

solving quadratic equations by factoring - 6x (2x-5) = 0 I need step by step breakdown and help...

OpenStudy (anonymous):

Solve each term in the product separately. Split into 2 equations. 6x=0 and 2x-5=0 So from the first equation: 6x=0, you get x=0 when you divide both sides by 6. And the second equation: 2x-5=0, carry the 5 across the equal sign to give you: 2x=5, then divide both sides by to get x=5/2

OpenStudy (anonymous):

so then X will be equal to 0 and 5/2

OpenStudy (anonymous):

I have a turd load of problems to do with this subject (solving quadratic equations by factoring) and I gotta be honest, I am struggling with it... if your gonna be on for a while, I could use the help.

OpenStudy (anonymous):

yes x will be equal to 0 and 5/2. Yes sure I'll be online for a while if you need any more help